Radioactive nuclei has10 ^8,its half life is 30s the number of nuclei left after 15s is

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
0.693/t1/2 =  
t1/2= half life 
=disintegration or rot consistent 
=>0.693/30 =  
=>=0.0231 
=2.303log(a/a-x)/t 
a=initial sum 
a-x= sum left after deterioration 
t=time of rot 
=>0.0231=2.303log(108/108-x)/15 
=>108/108-x=1.41 
=>Amount left=108-x=7,09,21,986 nuclei.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)